OpenStudy (anonymous):

which of the following possible base values of k is not consistent if y=f(x) is continuous, monotonic, and concave down for all values of x? a. 0 b. -2 c. 0.5 d. 6 e. 1

OpenStudy (anonymous):

0, because there would not able a change in y value

OpenStudy (anonymous):

can you explain further? sorry i'm so confused

OpenStudy (anonymous):

if the function is concave down, it means that f''(x)<0.

OpenStudy (anonymous):

by the way, what kind of funciton is this? exponential?

OpenStudy (anonymous):

it isn't specified

OpenStudy (anonymous):

If your teacher doesn't put in the effort to write good questions, then don't put in the effort to answer them.

OpenStudy (anonymous):

actually, I think it's -2, because if we assume it's an exponential funcion, y=-2^x is only one that's concave down

OpenStudy (anonymous):

what do you interpret base values of k to mean

OpenStudy (anonymous):

what do you mean K to mean?

OpenStudy (anonymous):

@lgbasallote @TuringTest @jim_thompson5910 , I need some help here

OpenStudy (anonymous):

this question is too vague :/

jimthompson5910 (jim_thompson5910):

If y = f(x) is continuous, monotonic and concave down, this means that f(x) is always decreasing over the interval (-infinity, infinity) Now if f(x) = k*g(x), then the above is only true if k is nonzero (g(x) is also continuous, monotonic and concave down). This my guess as to what/where k is if k were zero, then f(x) = 0, which would mean that it wouldn't be concave down anymore

OpenStudy (anonymous):

so it would be at 0 because that's where it would no longer be concave down

jimthompson5910 (jim_thompson5910):

that's my guess anyway, yes

OpenStudy (anonymous):

thanks so much for your help but it was -2 :(

jimthompson5910 (jim_thompson5910):

hmm then I'm thinking of something completely different

jimthompson5910 (jim_thompson5910):

oh, if you had f(x) = k^x or something like that and k was negative, then it wouldn't be continuous for all real values of x. So that's why k = -2 doesn't work

jimthompson5910 (jim_thompson5910):

but of course, that's all based on the assumption that f(x) = k^x or something similar to it